6. The following data are compiled during water of hydration experiment of BaCl2.xH2O.

Mass of beaker + watch glass 35.82 grams
Mass of beaker + watch glass + sample 37.80 grams
Mass of beaker + watch glass + residue after first heating 37.52 grams

Mass of beaker + watch glass + residue after second heating 37.51 grams

What is the formula of the hydrate?
Answer #1

(Mass of beaker + watch glass + sample 37.80 grams) - (Mass of beaker + watch glass 35.82 grams) = Mass of sample

= 1.98 grams

Mass of beaker + watch glass + residue after first heating 37.52 grams

Mass of water escaped = 37.8 - 37.52 = 0.28 grams

Number of moles of water escaped = 0.28/18 =0.0155

Mass of BaCl2 left = 1.98-0.28 = 1.7 grams

Moles of BaCl2 = 1.7g / 208.23 g/mol = 0.00861

Number of moles of H2O / Moles of BaCl2 = 0.0155 / 0.0861 = 1.8

Which is approximately equals to 2

Formula of the hydrate = BaCl2. 2H2O